OpenStudy (anonymous):

Which of the following is a recursive formula for the sequence whose first four terms are 4,2,-4, and -22? A. a^n = 3a_(n-1)-1 B. a^n=3a_(n-1)-10 C. a^n=-3a_(n-1)+1 D. a^n=-3a_(n-1)-10 2. Find the first six terms of the sequence in which a_1=4 and a_(n+1)+(n+6), for n≥1. A. 4,11,19,28,38,49 B. 4,10,16,22,28,34 C. 4,12,21,31,42,54 D. 4,10,18,27,37,48

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

2= 3(4) -10 -4=3(2) -10 -22= 3(-4)-10 What is the answer?

OpenStudy (anonymous):

This was for the first one

OpenStudy (anonymous):

The second question is not written correctly

OpenStudy (anonymous):

B. for the first one?

OpenStudy (anonymous):

The second one is written: a_1=4 and a_n+1 = a_n + (n+6), for n≥1

OpenStudy (anonymous):

for the first question is B the correct answer?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

B for the first one

OpenStudy (anonymous):

4 11 = 4 +6+1 19= 11 + 7+1 28= 19 +8+1 39 =28 +9+1 50= 39 + 10+1 None of the choices are right

OpenStudy (anonymous):

I got b, but i am not confident about my answer.

OpenStudy (anonymous):

{4,11,19,28,39,50} is not one of the choices. There is something wrong with the choices. Did you type them right:

OpenStudy (anonymous):

4=a_1 a_2 = a_1 + 6+1=4+7 =11

OpenStudy (anonymous):

The second term should be 11. I am going to gym

OpenStudy (anonymous):

oh ok . thank you for your help :)